# CMSO-transducing tree-like graph decompositions

### v2

@@ 5 Transducing split and bi-join decompositions
5.1 Transducing weakly-bipartitive trees
@@
To model bipartition systems, we use the extended vocabulary $\{\mathsf{BIPART}\}$ where $\mathsf{BIPART}$ is a set predicate name of arity $2$. A bipartition system $(U,\mathcal{B})$ is thus naturally modeled as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$ with universe $U_{\mathbb{B}}=U$ and interpretation $\mathsf{BIPART}_{\mathbb{B}}=\mathcal{B}$.
Two bipartitions $\{X,Y\}$ and $\{X^{\prime},Y^{\prime}\}$ of a set $U$ overlap if the four sets $X\cap X^{\prime}$, $X\cap Y^{\prime}$, $Y\cap X^{\prime}$ and $Y\cap Y^{\prime}$ are non-empty. A bipartition family is said to be laminar (aka overlap-free) if no two bipartitions in $\mathcal{B}$ overlap. By extension, we call a family $\mathcal{B}$ of bipartition of a set $U$ laminar whenever $(U,\mathcal{B})$ is a bipartition system which is laminar. Similarly to laminar set systems, we can associate a tree with each laminar bipartition system. The tree $T$, called the laminar tree induced by $(U,\mathcal{B})$ (or laminar tree of $(U,\mathcal{B})$), associated with a laminar bipartition system $(U,\mathcal{B})$ is unrooted, each element of $U$ corresponds to a leaf of $T$ and for each bipartition $\{X,Y\}$ in $\mathcal{B}$ there is exactly one edge $e$ such that $X=L(T_{1})$ and $Y=L(T_{2})$ for the two connected components $T_{1},T_{2}$ of $T-e$. We remark that there is a unique such tree and each inner node has degree at least three. Furthermore, the size of the laminar tree is linearly bounded in the size of the universe $U$.
A bipartition system $(U,\mathcal{B})$ is said to be weakly-bipartitive if for every two overlapping bipartitions $\{X,Y\},\{X^{\prime},Y^{\prime}\}\in\mathcal{B}$, the biparitions $\{X\cup X^{\prime},Y\cap Y^{\prime}\}$, $\{X\cup Y^{\prime},Y\cap X^{\prime}\}$, $\{Y\cup X^{\prime},X\cap Y^{\prime}\}$ and $\{Y\cup Y^{\prime},X\cap X^{\prime}\}$ are in $\mathcal{B}$. It is bipartitive, if additionally for every two overlapping bipartitions $\{X,Y\},\{X^{\prime},Y^{\prime}\}\in\mathcal{B}$, the bipartition $\{X\operatorname*{\bigtriangleup}X^{\prime},Y\operatorname*{\bigtriangleup}Y^{\prime}\}$ is also in $\mathcal{B}$. By extension, we call a family $\mathcal{B}$ of bipartitions of a set $U$ weakly-bipartitive or bipartitive whenever $(U,\mathcal{B})$ is a bipartition system which is weakly-bipartitive or partitive, respectively.
@@
The following lemma gives a connection between laminar trees of set systems and laminar trees of bipartition systems. Note that a very similar statement follows from [25, Lemma 1.14] and hence we only provide a proof for sake of completeness in the appendix.
Lemma 5.1.
@@
Lemma 5.2.
@@
Let $\Sigma$ be an extended vocabulary, including a set unary predicate name $\mathsf{BIPART}$ and not including the binary relational symbol $\mathsf{t}\text{-}\mathsf{edge}$. There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each bipartition system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$ and inducing the laminar tree $T$ with ${L(T)=U}$, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{B}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{T}$ representing $T$.
@@
Let $\mathcal{B}$ be a family of bipartitions of $U$ represented as the extended relational structure $\mathbb{B}$ and $\mathbb{A}$ a $\Sigma$-structure with $\mathbb{B}\sqsubseteq\mathbb{A}$. We first use the following two simple atomic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction to obtain a set system from $(U,\mathcal{B})$. Guess any colouring $A$ for which there is a single $a\in U$ with $A(a)$ being satisfied. Interpret using the following formula to obtain a $\mathsf{SET}$-predicate:
$\phi_{\mathsf{SET}}(X):=\forall a(A(a)\rightarrow(X=\{a\}\lor X=U\lor a\notin X))\land\mathsf{BIPART}(X,U\setminus X).$
The resulting set system $(U,\mathcal{S}_{a})$ represented by the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ is a set system by Lemma 5.1. Hence, we can use the transduction from Lemma 4.1 to obtain the $\Sigma\cup\{\mathsf{ancestor}\}$-structure $\mathcal{A}\sqcup\mathcal{T}$ where $\mathcal{T}$ is the $\{\mathsf{ancestor}\}$-structure representing the laminar tree $T$ of $(U,\mathcal{S}_{a})$. We now use the following two atomic transductions to obtain the weakly-bipartitive tree of $(U,\mathcal{B})$ according to Lemma 5.1. We fist restrict the universe to all nodes which are not the root of $\mathbb{T}$. We then interpret using the following formula to obtain the $\mathsf{t}\text{-}\mathsf{edge}$-predicate which makes the graph undirected and adds an edge from $a$ to the node representing the set $U\setminus\{a\}$ which is the root of $T$ after removing the node corresponding to $U$:
$\phi_{\mathsf{t}}\text{-}\mathsf{edge}(x,y):=\mathsf{parent}(x,y)\lor\mathsf{parent}(y,x)\lor(x=a\land\lnot\exists z\mathsf{parent}(z,y))\lor(y=a\land\lnot\exists z\mathsf{parent}(z,x)).$
Finally, the transduction keeps only $\mathsf{t}\text{-}\mathsf{edge}$ and any $Q\in\Sigma$ as relations for its output.
Similar to the set system case, for each weakly-bipartitive family $\mathcal{B}$ we can equip the laminar tree with a labelling of its inner nodes and a partial order which characterizes the family $\mathcal{B}$.
Theorem 5.4 (Theorem 3 in [17]).
Let $\mathcal{B}$ be a weakly-bipartitive family, $\mathcal{B}_{!}$ be its subfamily of strong bipartitions, and $T$ the laminar tree it induces. There exists a total labeling function $\lambda$ from the set $V(T)\setminus L(T)$ of the inner nodes of $T$ to the set $\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a linear ordering $<_{t}$ of its neighbors, such that every inner node of degree exactly three is labeled by $\mathsf{degenerate}$ and the following conditions are satisfied:
  • for each bipartition $\{X,Y\}\in\mathcal{B}\setminus\mathcal{B}_{!}$, there exists $t\in V(T)\setminus L(T)$ and a subset $\mathcal{C}$ of neighbors of $t$ such that either $X$ or $Y$ is equal to the set $\bigcup_{c\in\mathcal{C}}L(T^{t}_{c})$ and either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $<_{t}$-interval, or $\lambda(t)=\mathsf{degenerate}$;
  • conversely, for each inner node $t$ and each non-empty subset $\mathcal{C}$, $\mathcal{C}\not=U$ of neighbors of $t$, the bipartition $\big{\{}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c}),\bigcup_{c\notin\mathcal{C}}L(T^{t}_{c})\big{\}}$ is a member of $\mathcal{B}$ if either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is $<_{t}$-interval or $\lambda(t)=\mathsf{degenerate}$.
The tree $T$ and its labelling function $\lambda$ are uniquely determined by $\mathcal{B}$. We note that the total orders $<_{t}$ are uniquely determined by $\mathcal{B}$ up to inverting and cyclic shifting, i.e., for a linear order $<$ of $X=\{x_{1},\dots,x_{\ell}\}$ with $x_{1}<\dots<x_{\ell}$ we call the order $<^{\prime}$ of $X$ with $x_{\ell}<^{\prime}x_{1}<^{\prime}\dots<^{\prime}x_{\ell-1}$ a cyclic shift of $<$. To see this, note that in the theorem above the bipartitions obtained from nodes with label $\mathsf{linear}$ can be equivalently obtained from a $<_{t}$-interval or the complement of a $<_{t}$-interval. Indeed, a cyclic shifting of a linear order $<$ maintains the property of being either a $<$-interval or the complement of a $<$-interval. Neglecting that $<_{t}$ are only unique up to inverting and cyclic shifting, we represent every weakly-bipartitive family $\mathcal{B}$ by a triple $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ where $T$ is the laminar tree induced by $\mathcal{B}_{!}$, $\lambda:V(T)\setminus L(T)\to\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$ is the labeling function, and $<_{t}$ is the total order of the neighbours of $t$ described above. Note that in this case $\bigcup_{t\in\lambda^{-1}(\mathsf{linear})}<_{t}$ is not a partial order as it is not necessarily transitive. We call the triple $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ from Theorem 5.4 the bipartitive tree of $\mathcal{B}$.
We model weakly-bipartitive trees $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ of a weakly-bipartitive family $\mathcal{B}$ of bipartition of $U$ by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate},\mathsf{betweeness}\}$-structure $\mathbb{T}$ of universe $U_{\mathbb{T}}=V(T)$ such that $\langle V(T),\mathsf{t}\text{-}\mathsf{edge}_{\mathbb{T}}\rangle\sqsubseteq\mathbb{T}$ models $T$ with ${L(T)=U}$, $\mathsf{degenerate}_{\mathbb{T}}$ is a unary relation which selects all inner nodes of $T$ of label $\mathsf{degenerate}$, i.e., ${\mathsf{degenerate}_{\mathbb{T}}=\lambda^{-1}(\mathsf{degenerate})}$, and $\mathsf{betweeness}_{\mathbb{T}}$ is a relation of arity four selecting all quadruples $(t,x,y,z)$ for which $y$ is between $x$ and $z$ in $<_{t}$. Note that in order to define $\mathsf{betweeness}$ we essentially choose one representative order amongst all orders obtained by inverting and cyclic shifting. Similarly to the weakly-partitive case, we can recover the inner nodes of $T$ that are labeled by $\mathsf{linear}$ through an $\operatorname{MSO}$-formula as those inner nodes whose neighbors are related by $\mathsf{betweeness}$. Then $\mathsf{prime}$ nodes are the ones which are labeled neither by $\mathsf{degenerate}$ nor by $\mathsf{linear}$ which can be obtained by an $\operatorname{MSO}$-transduction. Furthermore, Theorem 5.4, gives rise to an $\operatorname{MSO}$-transduction which takes as input a weakly-bipartitive tree and outputs the weakly-bipartitive set system which induced it.
We can now extend the transduction from Lemma 5.2 to obtain a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction producing the weakly-bipartitive tree of a weakly-bipartitive family.
Theorem 5.5.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for every weakly-bipartitive bipartition system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $\tau(\mathbb{B})$ is non-empty and every output in $\tau(\mathbb{B})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate},\mathsf{betweeness}\}$-structure $\mathbb{T}$ representing the weakly-bipartitive tree $\left(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})}\right)$ of $(U,\mathcal{B})$.
Proof 5.6.
@@
Lemma 5.2 already provides a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which on input $\mathbb{B}$ outputs the $\{\mathsf{BIPART},\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{B}\sqcup\mathbb{T}^{\prime}$, where $\mathbb{T}^{\prime}$ represents the $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure modeling the laminar tree $T$ of $(U,\mathcal{B})$. To complete the proof of the theorem, we provide an $\operatorname{MSO}$-transduction which outputs $\mathbb{T}$ given input $\mathbb{B}\sqcup\mathbb{T}^{\prime}$. For the transduction we define an $\operatorname{MSO}$-formula $\phi_{\mathsf{degenerate}}$ which selects the inner nodes of $T$ which are labeled $\mathsf{degenerate}$, and a formula $\phi_{\mathsf{betweeness}}$ which selects quadruples of nodes $(t,x,y,z)$ such that either $x<_{t}y<_{t}z$ or $z<_{t}y<_{t}x$. Recall, that the ordering $<_{t}$ for each inner node $t$ of the neighbours of $t$ are only unique up to inverting the order and cyclic shift. In order to select a representative of $<_{t}$ which is unique only up to inverting the order, we need non-determinism (details follow). We use a functional symbol $\mathsf{leafset}(t,s)$ which is interpreted as the set $L(T_{s}^{t})$. The functional predicate $\mathsf{leafset}$ clearly can be defined in $\operatorname{MSO}$ as a element $z$ belongs to $\mathsf{leafset}(t,s)$ if and only if $z$ is a leaf and the path from $z$ to $s$ does not contain $t$.
@@
Observe that by Theorem 5.4 an inner node $t$ is labeled $\mathsf{degenerate}$ if and only if for every pair of neighbours $s_{1},s_{2}$ of $t$ the bipartition $\big{\{}L(T^{t}_{s_{1}})\cup L(T^{t}_{s_{2}}),U\setminus\big{(}L(T^{t}_{s_{1}})\cup L(T^{t}_{s_{2}})\big{)}\big{\}}$ is in $\mathcal{B}$. We can define this in $\operatorname{MSO}$ using the following formula:

### v4

@@ 5. Transducing split and bi-join decompositions
5.1. Transducing weakly-bipartitive trees
@@
To model bipartition systems, we use the extended vocabulary $\{\mathsf{BIPART}\}$ where $\mathsf{BIPART}$ is a unary set predicate name. A bipartition system $(U,\mathcal{B})$ is thus naturally modeled as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$ with universe $U_{\mathbb{B}}=U$ and interpretation $\{\{X,U\setminus X\}\mid X\in\mathsf{BIPART}_{\mathbb{B}}\}=\mathcal{B}$.
Two bipartitions $\{X,Y\}$ and $\{X^{\prime},Y^{\prime}\}$ of a set $U$ overlap if the four sets $X\cap X^{\prime}$, $X\cap Y^{\prime}$, $Y\cap X^{\prime}$ and $Y\cap Y^{\prime}$ are non-empty. A bipartition family is said to be laminar (aka overlap-free) if no two bipartitions in $\mathcal{B}$ overlap. By extension, we call a family $\mathcal{B}$ of bipartition of a set $U$ laminar whenever $(U,\mathcal{B})$ is a bipartition system which is laminar. Similarly to laminar set systems, we can associate a tree with each laminar bipartition system. The tree $T$, called the laminar tree induced by $(U,\mathcal{B})$ (or laminar tree of $(U,\mathcal{B})$), associated with a laminar bipartition system $(U,\mathcal{B})$ is unrooted, each element of $U$ corresponds to a leaf of $T$ and for each bipartition $\{X,Y\}$ in $\mathcal{B}$ there is exactly one edge $e$ such that $X=L(T_{1})$ and $Y=L(T_{2})$ for the two connected components $T_{1},T_{2}$ of TeT-e. We remark that there is a unique such tree and each inner node has degree at least three. Furthermore, the size of the laminar tree is linearly bounded in the size of the universe $U$.
A bipartition system $(U,\mathcal{B})$ is said to be weakly-bipartitive if for every two overlapping bipartitions $\{X,Y\},\{X^{\prime},Y^{\prime}\}\in\mathcal{B}$, the biparitions $\{X\cup X^{\prime},Y\cap Y^{\prime}\}$, $\{X\cup Y^{\prime},Y\cap X^{\prime}\}$, $\{Y\cup X^{\prime},X\cap Y^{\prime}\}$ and $\{Y\cup Y^{\prime},X\cap X^{\prime}\}$ are in $\mathcal{B}$. It is bipartitive, if additionally for every two overlapping bipartitions $\{X,Y\},\{X^{\prime},Y^{\prime}\}\in\mathcal{B}$, the bipartition $\{X\operatorname*{\bigtriangleup}X^{\prime},Y\operatorname*{\bigtriangleup}X^{\prime}\}$ is also in $\mathcal{B}$. By extension, we call a family $\mathcal{B}$ of bipartitions of a set $U$ weakly-bipartitive or bipartitive whenever $(U,\mathcal{B})$ is a bipartition system which is weakly-bipartitive or partitive, respectively.
Figure 7. Bipartitions to be included in a weakly-bipartitive system of bipartitions (excluds the rightmost) or in a bipartitive system bipartitions (includes the rightmost) for two overlapping bipartitions $\{X,Y\}$ and $\{X^{\prime},Y^{\prime}\}$.
@@
The following lemma gives a connection between laminar trees of set systems and laminar trees of bipartition systems. Note that a very similar statement follows from [Rao06, Lemma 1.14] and hence we only provide a proof for sake of completeness.
Lemma 20.
@@
Proof 5.1.
To argue the second assertion, assume that $T$ is the laminar tree of the set system $(U,\mathcal{S}_{a})$. First observe that in the set system $(U,\mathcal{S}_{a})$ the element $a$ is not contained in any set apart from $U$ and $\{a\}$ and hence the leaf $\{a\}$ is a child of the root $U$ of $T$. Additionally, the set $U\setminus\{a\}$ is in $S_{a}$ as the bipartition $(\{a\},U\setminus\{a\})$ is in $\mathcal{B}$ by definition and hence $U\setminus\{a\}$ is also a child of the root $U$ of $T$. Let $T^{\prime}$ be the tree obtained from $T$ by removing node $U$, making all edges undirected and adding an edge from node $\{a\}$ to node $U\setminus\{a\}$. Indeed, by our previous observations, $T^{\prime}$ is an undirected tree. Furthermore, there is a correspondence between elements of $U$ and the leaves of $T^{\prime}$ (element $b\in U$ corresponds to leaf $\{b\}$ of $T^{\prime}$). To check that $T^{\prime}$ is indeed the laminar tree of $T$, we need to verify that for each bipartition $\{X,Y\}$ in $\mathcal{B}$ there is exactly one edge $e$ of $T^{\prime}$ such that $X=L(T^{\prime}_{1})$ and $Y=L(T^{\prime}_{2})$ for the two connected components $T^{\prime}_{1},T^{\prime}_{2}$ of $T^{\prime}-e$. We say that the edge $e$ implements the bipartition $\{X,Y\}$. Assume that $\{X,Y\}$ is any bipartition from $\mathcal{B}$. First consider the case that $\{X,Y\}$ is the bipartition $\{\{a\},U\setminus\{a\}\}$. Then $\{X,Y\}$ is clearly implemented by the edge $\{\{a\},U\setminus\{a\}\}$ of $T^{\prime}$. Now assume that $\{X,Y\}$ is any other bipartition in $\mathcal{B}$ and assume, without loss of generality, that $a\in Y$. Hence, $X\in\mathcal{S}_{a}$. Let $X^{\prime}$ be the parent of $X$ in $T$. Note that $X$ must be a proper subset of $U\setminus\{a\}$ as by our assumption $\{X,Y\}$ is not the partition $\{\{a\},U\setminus\{a\}$ and hence $X^{\prime}$ cannot be $U$. As we only deleted edge incident to $U$ in our construction of $T^{\prime}$, $\{X,X^{\prime}\}$ is an edge in $T^{\prime}$ which clearly implements the bipartition $\{X,Y\}$. Finally, every bipartition $\{X,Y\}\in\mathcal{B}$ is clearly implemented by only one edge as every inner node of the tree $T^{\prime}$ has degree at least three. Therefore, $T^{\prime}$ is the laminar tree of $(U,\mathcal{B})$.
Lemma 21.
@@
Let $\Sigma$ be an extended vocabulary, including a unary set predicate name $\mathsf{BIPART}$ and not including the binary relational symbol $\mathsf{t}\text{-}\mathsf{edge}$. There exists a non-deterministic overlay $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each bipartition system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$ and inducing the laminar tree $T$ with L(T)=U{L(T)=U}, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{B}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to $\mathbb{A}\sqcup\mathbb{T}$ for some $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{T}$ representing $T$.
Proof 5.2.
Let $\mathcal{B}$ be a family of bipartitions of $U$ represented as the extended relational structure $\mathbb{B}$ and $\mathbb{A}$ a $\Sigma$-structure with $\mathbb{B}\sqsubseteq\mathbb{A}$. We first use the following two simple atomic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions to obtain a set system from $(U,\mathcal{B})$. Guess any colouring $A$ for which there is a single $a\in U$ with $A(a)$ being satisfied. Interpret using the following formula to obtain a $\mathsf{SET}$-predicate:
$\phi_{\mathsf{SET}}(X):=\forall a(A(a)\rightarrow(X=\{a\}\lor X=U\lor a\notin X))\land(\mathsf{BIPART}(X)\lor\mathsf{BIPART}(U\setminus X)).$
The resulting set system $(U,\mathcal{S}_{a})$ represented by the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ is a set system by 20. Hence, we can use the transduction from 14 to obtain the $\Sigma\cup\{\mathsf{ancestor}\}$-structure $\mathbb{A}\sqcup\mathbb{T}$ where $\mathbb{T}$ is the $\{\mathsf{ancestor}\}$-structure representing the laminar tree $T$ of $(U,\mathcal{S}_{a})$. We now use the following two atomic transductions to obtain the weakly-bipartitive tree of $(U,\mathcal{B})$ according to 20. We first restrict the universe to all nodes which are not the root of $\mathbb{T}$. We then interpret using the following formula to obtain the $\mathsf{t}\text{-}\mathsf{edge}$-predicate which makes the graph undirected and adds an edge from $a$ to the node representing the set $U\setminus\{a\}$ which is the root of $T$ after removing the node corresponding to $U$:
$\phi_{\mathsf{t}}\text{-}\mathsf{edge}(x,y):=$ $\mathsf{parent}(x,y)\lor\mathsf{parent}(y,x)\lor$
$(x=a\land y\not=a\land\lnot\exists z\mathsf{parent}(z,y))\lor(x\not=a\land y=a\land\lnot\exists z\mathsf{parent}(z,x)).$
Finally, the transduction keeps only $\mathsf{t}\text{-}\mathsf{edge}$ and any $Q\in\Sigma$ as relations for its output. We do not need to filter the outputs as any guess for $A$ yields a valid laminar tree.
Similar to the set system case, for each weakly-bipartitive family $\mathcal{B}$ we can equip the laminar tree with a labelling of its inner nodes and a partial order which characterizes the family $\mathcal{B}$. We remind the reader that for a node $t$ of a tree $T$ with neighbour $s$ we denote by $T^{t}_{s}$ the connected component of TtT-t containing $s$.
{thmC}
[[dM03, Theorem 3]] Let $\mathcal{B}$ be a weakly-bipartitive family, $\mathcal{B}_{!}$ be its subfamily of strong bipartitions, and $T$ the laminar tree it induces. There exists a total labeling function $\lambda$ from the set $V(T)\setminus L(T)$ of the inner nodes of $T$ to the set $\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a linear ordering $_{t}$ of its neighbours, such that every inner node of degree exactly three is labeled by $\mathsf{degenerate}$ and the following conditions are satisfied:
  • for each bipartition $\{X,Y\}\in\mathcal{B}\setminus\mathcal{B}_{!}$, there exists $t\in V(T)\setminus L(T)$ and a subset $\mathcal{C}$ of neighbours of $t$ such that either $X$ or $Y$ is equal to the set $\bigcup_{c\in\mathcal{C}}L(T^{t}_{c})$ and either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $_{t}-$interval, or $\lambda(t)=\mathsf{degenerate}$;
  • conversely, for each inner node $t$ and each non-empty subset $\mathcal{C}$, $\mathcal{C}\not=U$ of neighbours of $t$, the bipartition $\big{\{}\bigcup_{c\in\mathcal{C}}L(T^{t}_{c}),\bigcup_{c\notin\mathcal{C}}L(T^{t}_{c})\big{\}}$ is a member of $\mathcal{B}$ if either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is $_{t}-$interval or $\lambda(t)=\mathsf{degenerate}$.
The tree $T$ and its labelling function $\lambda$ are uniquely determined by $\mathcal{B}$. We note that the total orders $_{t}$ are uniquely determined by $\mathcal{B}$ up to inverting and cyclic shifting, i.e., for a linear order << of $X=\{x_{1},\dots,x_{\ell}\}$ with $x_{1}<\dots<x_{\ell}$ we call the order $^{\prime}$ of $X$ with $x_{\ell}<^{\prime}x_{1}<^{\prime}\dots<^{\prime}x_{\ell-1}$ a cyclic shift of <<. To see this, note that in the theorem above the bipartitions obtained from nodes with label $\mathsf{linear}$ can be equivalently obtained from a $_{t}-$interval or the complement of a $_{t}-$interval. Indeed, a cyclic shifting of a linear order << maintains the property of being either a <<-interval or the complement of a <<-interval. Neglecting that $_{t}$ are only unique up to inverting and cyclic shifting, we represent every weakly-bipartitive family $\mathcal{B}$ by a triple $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ where $T$ is the laminar tree induced by $\mathcal{B}_{!}$, $\lambda:V(T)\setminus L(T)\to\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$ is the labeling function, and $_{t}$ is the total order of the neighbours of $t$ described above. Note that in this case $\bigcup_{t\in\lambda^{-1}(\mathsf{linear})}<_{t}$ is not a partial order as it is not necessarily transitive. We call the triple $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ from 5.2 the bipartitive tree of $\mathcal{B}$.
We model weakly-bipartitive trees $(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})})$ of a weakly-bipartitive family $\mathcal{B}$ of bipartition of $U$ by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate},\mathsf{cross}\}$-structure $\mathbb{T}$ of universe $U_{\mathbb{T}}=V(T)$ such that $\langle V(T),\mathsf{t}\text{-}\mathsf{edge}_{\mathbb{T}}\rangle\sqsubseteq\mathbb{T}$ models $T$ with L(T)=U{L(T)=U}, $\mathsf{degenerate}_{\mathbb{T}}$ is a unary relation which selects all inner nodes of $T$ of label $\mathsf{degenerate}$, i.e., $\mathsf{degenerate}_{\mathbb{T}}=\lambda^{-1}(\mathsf{degenerate})$}, and $\mathsf{cross}_{\mathbb{T}}$ is a relation of arity five selecting all 5-tuples $(t,w,x,y,z)$ for which the chord $wy$ crosses the chord $xz$ in $_{t}$ understood as a cyclic order. Similarly to the weakly-partitive case, we can recover the inner nodes of $T$ that are labeled by $\mathsf{linear}$ through an $\operatorname{MSO}$-formula as those inner nodes whose neighbours are related by $\mathsf{cross}$. Then $\mathsf{prime}$ nodes are the ones which are labeled neither by $\mathsf{degenerate}$ nor by $\mathsf{linear}$ which can be obtained by an $\operatorname{MSO}$-transduction. Furthermore, 5.2, gives rise to an $\operatorname{MSO}$-transduction which takes as input a weakly-bipartitive tree and outputs the weakly-bipartitive set system which induced it.
We can now extend the transduction from 21 to obtain a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction producing the weakly-bipartitive tree of a weakly-bipartitive family.
Theorem 22.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for every weakly-bipartitive bipartition system $(U,\mathcal{B})$ represented as the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $\tau(\mathbb{B})$ is non-empty and every output in $\tau(\mathbb{B})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate},\mathsf{cross}\}$-structure $\mathbb{T}$ representing the weakly-bipartitive tree $\left(T,\lambda,(<_{t})_{t\in\lambda^{-1}(\mathsf{linear})}\right)$ of $(U,\mathcal{B})$.
@@
21 already provides a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which on input $\mathbb{B}$ outputs the $\{\mathsf{BIPART},\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{B}\sqcup\mathbb{T}^{\prime}$, where $\mathbb{T}^{\prime}$ represents the $\{\mathsf{t}\text{-}\mathsf{edge}\}$-structure modeling the laminar tree $T$ of $(U,\mathcal{B})$. To complete the proof of the theorem, we provide an $\operatorname{MSO}$-transduction which outputs $\mathbb{T}$ given input $\mathbb{B}\sqcup\mathbb{T}^{\prime}$. For the transduction we define an $\operatorname{MSO}$-formula $\phi_{\mathsf{degenerate}}$ which selects the inner nodes of $T$ which are labeled $\mathsf{degenerate}$, and a formula $\phi_{\mathsf{cross}}$ which selects 5-tuples of nodes $(t,w,x,y,z)$ such that either $w<_{t}x<_{t}y<_{t}z$ or $z<_{t}y<_{t}x<_{t}w$ or some cyclic shift of this. Recall, that the ordering $_{t}$ for each inner node $t$ of the neighbours of $t$ are only unique up to inverting the order and cyclic shift. We use a functional symbol $\mathsf{leafset}(t,s)$ which is interpreted as the set $L(T_{s}^{t})$ where $T_{s}^{t}$ is the connected component of TtT-t containing $s$. The functional predicate $\mathsf{leafset}$ clearly can be defined in $\operatorname{MSO}$ as an element $z$ belongs to $\mathsf{leafset}(t,s)$ if and only if $z$ is a leaf and the path from $z$ to $s$ does not contain $t$.
@@
Observe that by 5.2 an inner node $t$ is labeled $\mathsf{degenerate}$ if and only if for every pair of neighbours $s_{1},s_{2}$ of $t$ the bipartition $\big{\{}L(T^{t}_{s_{1}})\cup L(T^{t}_{s_{2}}),U\setminus\big{(}L(T^{t}_{s_{1}})\cup L(T^{t}_{s_{2}})\big{)}\big{\}}$ is in $\mathcal{B}$. We can define this in $\operatorname{MSO}$ using the following formula: